OpenStudy (anonymous):

A line passes through the point (-4,3) and has a slope of 5/2 Write an equation in slope-intercept form for this line.

OpenStudy (anonymous):

you have a pt and a slope... use pt slope form. y-y1 = m(x-x1) so y-3=5/2(x+4)

OpenStudy (anonymous):

now distribute the 5/2. and then add the three. and you will have your answer in slope intercept form!

OpenStudy (anonymous):

y = (5 x)/2+13

OpenStudy (anonymous):

Looks good!